OpenStudy (anonymous):

can anyone tell me if I worked this problem right and got the right answer. I got 2 answers though. the problem is through (2,-4) and (0,-7) find slope intercept, I got y=3/2x-7 and y=6/7x-7. did I do the problem right

OpenStudy (anonymous):

@Zelda when u get a chance please return to my question

OpenStudy (anonymous):

Well, let's go though it and check :D Slope = m = (-7-(-4))/0-2 = -7+4/-2 = -3/-2 = 3/2 We have m! Yay! now, we'll choose one of the two points, so the second one say (0,-7). Equation in point slope is: y=mx+b We have m,x,y solve for b So we do: -7 = (3/2) * 0 +b -7 = b. So we have b now too, yay :D Now just put it in final equation: \[y=\frac{ 3 }{ 2 }x-7\] So the first equation is right :D

OpenStudy (anonymous):

I am so glad. thank you so much. I am understanding these problems better and better. YAYYYYYY

OpenStudy (anonymous):

Yay for you too :D Hope this helped.

OpenStudy (anonymous):

YES IT DID THANK YOU

OpenStudy (anonymous):

No problem kiddo, mind marking it best response?

OpenStudy (anonymous):

I got ya